# Unleashing the Power of LLMs as Multi-Modal Encoders for Text and Graph-Structured Data

### v1

@@ 4.3.3. Results
BERT-Base (Devlin et al., 2019) 13.3 2.2
RoBERTa-Large (Liu et al., 2019) 43.3 20.4
E5-Small (Wang et al., 2022) 65.6 34.8
E5-Base (Wang et al., 2022) 73.1 36.4
E5-Large (Wang et al., 2022) 72.6 38.6
GTR-XXL (Ni et al., 2022b) 66.2 46.7
SGPT (Muennighoff, 2022) 74.7 37.2
E5-Mistral (Wang et al., 2023) 76.1 53.5
@@
QA-GNN (Yasunaga et al., 2021) 41.4 19.5
GreaseLM (Zhang et al., 2022) 48.9 29.3
Janus (Ours) 82.9 $\pm$ 1.9 54.1 $\pm$ 0.1
As shown in Table 5, Janus achieves the best performance on both the SciFact and FiQA retrieval tasks. When comparing Janus with E5-Mistral, the backbone of our model and the best-performing baseline, we observe a clear improvement on both datasets. This performance boost highlights the effectiveness of incorporating graph-based contextual information into the retrieval process.
A closer examination of the baselines that integrate language models with knowledge graphs (LM + KG), such as QA-GNN and GreaseLM, reveals interesting trends. Both models use RoBERTa-Large to initialize node embeddings. From the results, GreaseLM outperforms the RoBERTa-Large model by a margin of 5.6% on SciFact and 8.9% on FiQA, which aligns with our earlier findings that graph context can be useful for enhancing retrieval performance. However, QA-GNN shows a performance degradation compared to RoBERTa-Large, particularly on FiQA (19.5 vs. 20.4). The reason for this drop may be attributed to QA-GNN’s use of a simple MLP for combining the graph and text embeddings. This weaker integration mechanism is likely insufficient for fully leveraging the graph context, resulting in suboptimal performance. In contrast, GreaseLM employs a shallow transformer, which offers better capacity for fusing multimodal information, leading to moderate gains.
Our method, Janus, demonstrates its superiority by using a powerful LLM backbone to combine the graph and text representations, which is reflected in the substantial performance improvements on both datasets. This highlights the advantage of using a more expressive model to combine graph and textual information.
4.4. Ablation Studies
To better understand the contributions of different components in Janus, we perform an ablation study on the CommonsenseQA and OpenBookQA datasets. The results, presented in Table 6, highlight the impact of key components on overall performance. Additionally, Figure 3 and Table 7 show the effects of different choices of components on model performance.
4.4.1. Effect of graph context information.
When Janus is input with only textual input, excluding graph context information, there is a significant drop in performance, with accuracy decreasing to 69.49% on CommonsenseQA and 74.80% on OpenBookQA. This sharp decline underscores the critical role that graph context plays in enhancing the model’s ability to understand and reason with the data. The graph provides structured knowledge that complements the unstructured text, and without it, the model relies solely on the textual input, which limits its capacity to effectively handle complex embedding tasks.
4.4.2. Effect of graph-text alignment.
Removing contrastive learning from Janus leads to a performance drop, with the accuracy decreasing from 81.09% to 79.69% on CommonsenseQA, and from 86.67% to 84.80% on OpenBookQA. This highlights the importance of contrastive learning in aligning graph and text embeddings, ensuring better integration of multimodal information.
We further investigate the effectiveness of graph-text alignment by studying how the alignment evolves during training. Specifically, we calculate the mean Euclidean distance between the normalized graph and text embeddings on the dev set (shown as the red dashed line) and compare it to the corresponding development accuracy (purple solid line) across different training epochs, as illustrated in Figure 3(c). This figure shows a clear inverse relationship between the two curves: as the graph-text distance decreases, accuracy improves. Notably, at epoch 2, the graph-text distance reaches its minimum, while the dev accuracy peaks. This trend suggests that better alignment between graph and text embeddings contributes directly to improved model performance, highlighting the effectiveness of our graph-text alignment strategy in Janus.
Table 6. Ablation study results on test accuracy.
Methods CommonsenseQA OpenBookQA
GreaseLM (Zhang et al., 2022) 74.20 $\pm$ 0.40 83.87 $\pm$ 1.29
Janus w/o graph 69.49 $\pm$ 0.28 74.80 $\pm$ 0.35
Janus w/o alignment 79.69 $\pm$ 0.85 84.80 $\pm$ 1.06
Janus 81.09 $\pm$ 0.73 86.67 $\pm$ 1.10
@@ 4.4. Ablation Studies
Figure 3. (a) The effect of LLM backbone choice on accuracy for the CommonsenseQA dataset. The figure shows three series: E5, LLaMA-3, and LLaMA-2, along with a single Mistral-7B model. (b) The effect of graph encoder depth (number of GNN layers) on test accuracy for CommonsenseQA and OpenBookQA datasets. The shaded areas represent the standard deviation, indicating the variance in performance across different trials. (c) Graph-text embedding distance (red dashed) and dev accuracy (purple solid) on CommonsenseQA across training epochs.
4.4.3. Effect of LLM Backbone Choice
@@
Janus exhibits flexibility in adopting various LLMs as its backbone. To evaluate the impact of different LLM backbones on performance, we compare four series of models: the E5 series (Wang et al., 2022, 2023), LLaMA-2 (Touvron et al., 2023), LLaMA-3 (Dubey et al., 2024), and Mistral (Jiang et al., 2023), as illustrated in Figure 3(a). We first observe that increasing the model size within each series consistently enhances the performance. This trend highlights that larger models, with more parameters, are typically able to better capture complex relationships in multi-modal data, leading to higher performance.
Additionally, LLaMA-3 significantly improves over LLaMA-2. Specifically, the performance of LLaMA-3-3B matches that of LLaMA-2-7B and LLaMA-3-8B achieves results comparable to LLaMA-2-13B, indicating that the new LLaMA-3 architecture brings considerable advancements compared to its predecessor LLaMA-2.
Furthermore, when comparing E5-Mistral to the base Mistral model, we see that E5-Mistral, which is a fine-tuned version of Mistral optimized for text embedding tasks, outperforms Mistral by a noticeable margin. This indicates that models specifically tailored for text embeddings are better suited for tasks involving joint embeddings of text and graph data.
4.4.4. Effect of Graph Encoder Depth
We evaluate the effect of varying the number of GAT layers in the graph encoder on the performance of Janus using the CommonsenseQA and OpenBookQA datasets. The results are shown in Figure 3(b). The darker lines in the figure represent the mean test accuracy, while the shaded regions around the lines indicate the standard deviation.
@@ 4.4.4. Effect of Graph Encoder Depth
As shown in Figure 3(b), both datasets exhibit slight fluctuations in test accuracy as the number of GAT layers increases from 3 to 7. For CommonsenseQA, the accuracy remains relatively stable across different GNN depths, fluctuating around 80%, with minor drops at 5 and 7 layers. This suggests that the method is quite robust to changes in GAT depth, with only small variations in performance. Similarly, for OpenBookQA, we also observe a stable trend, with accuracy consistently around 86%.
4.4.5. Impact of Node Embedding Initialization
Table 7. Impact of different graph node embedding initialization models on accuracy for the CommonsenseQA dataset.
Model Output Dimension Accuracy
RoBERTa-Large (Liu et al., 2019) 1,024 81.09 $\pm$ 0.73
E5-Small (Wang et al., 2022) 384 80.37 $\pm$ 0.38
E5-Base (Wang et al., 2022) 768 80.77 $\pm$ 0.05
E5-Large (Wang et al., 2022) 1,024 80.79 $\pm$ 0.30
E5-Mistral (Wang et al., 2023) 4,096 80.42 $\pm$ 0.90
Interestingly, even when using lighter models for node embedding initialization, such as E5-Small, which has a smaller output dimension (384), the performance remains competitive. This demonstrates that even when the node initialization model differs from the actual backbone used in Janus, there is no significant performance degradation. For example, compared with E5-Mistral node embedding initialization, the lighter E5-Small model offers the advantage of reduced computational cost without compromising much on accuracy.
5. Related Work
Graphs and Language Models for Multi-Modal Embedding Tasks. Integration of graph data with language models has been an evolving field, aiming to combine structured graph data with unstructured textual information for enhanced data analysis. Historically, early attempts in this area employed an MLP or a shallow transformer to merge the information from both modalities (Lin et al., 2019b; Yasunaga et al., 2021; Zhang et al., 2022; Feng et al., 2020), which may not fully exploit the rich contextual information from text and graph data.
The advent of LLMs has brought a transformative shift to the NLP field. LLMs, with their extensive pre-training on diverse corpora, offer unprecedented capabilities for deep semantic understanding and reasoining (Zhao et al., 2023a; Gao et al., 2023). Recent research has begun to explore the potential of LLMs for enhancing multimodal embedding tasks of graphs and text (Huang et al., 2024; Tian et al., 2024; Zhao et al., 2023b; He et al., 2024a; Zhu et al., 2024b). These studies primarily focus on augmenting the graph representation capabilities within the LLM framework. While these efforts mark significant advancements, they predominantly concentrate on graph representation learning rather than direct NLP tasks. Additionally, these methods do not explicitly align the semantic spaces of graphs and text. In contrast, our approach not only integrates LLMs for handling complex multimodal inputs but also introduces an explicit alignment mechanism between graph and text embeddings. This alignment is crucial for enhancing the semantic integration and boosting performance across diverse NLP tasks by capturing complex intermodal relationships.
There are recent studies that have explored using LLMs and graph data for generative tasks beyond embedding tasks (He et al., 2024b; Perozzi et al., 2024; Zhang et al., 2024; Chai et al., 2023; Hu et al., 2024; Jin et al., 2024). Our work, however, concentrates on embedding tasks, which is an orthogonal direction. Unlike generative tasks, which primarily focus on creating new content via next-token prediction, embedding tasks are aimed at developing rich, informative representations that can be used directly to enhance performance in downstream applications such as classification and retrieval.
Contrastive Learning. Contrastive learning has gained significant attention in recent years as a method to learn representations by maximizing agreement between positive pairs while pushing negative pairs apart in the embedding space (He et al., 2020; Chen et al., 2020; Lin et al., [n. d.], 2024; Yu et al., 2022). Pioneering works like SimCLR (Chen et al., 2020) and MoCo (He et al., 2020) have demonstrated the effectiveness of contrastive learning in the visual domain. Similar approaches have been applied to graphs (Yu et al., 2022; You et al., 2020; Fang et al., 2024) and language models (Iter et al., 2020b; Fang et al., 2020; Giorgi et al., 2021), but these methods typically operate within a single modality (e.g., graphs or text). In contrast, our approach introduces contrastive learning to align the graph embedding space with the text embedding space. By transforming graphs into textual descriptions and applying contrastive learning, we ensure that embeddings from both modalities align closely, enhancing the model’s ability to perform tasks that require both graph-based reasoning and text comprehension.
6. Conclusion
@@
In this paper, we have introduced Janus, a simple yet effective framework designed to integrate graph and text data using a novel alignment strategy and LLMs. We have demonstrated that Janus enhances the semantic coherence between these two modalities, resulting in significantly improved performance on several NLP tasks and datasets. By aligning graph embeddings directly with text embeddings, Janus ensures a deeper integration of structured and unstructured data. Future directions include exploring additional ways to fine-tune the alignment mechanisms to further enhance model performance, and investigate the scalability of our approach in handling larger datasets.
@@ Appendix A Ethical Use of Data and Informed Consent
Appendix B Dataset Details
B.1. KG-Contextulized QA
@@
CommonsenseQA (Talmor et al., 2019) is a 5-way multiple-choice QA dataset focused on applying commonsense knowledge in answering questions. It includes 12,102 questions, each with one correct answer and four distractor answers.
@@
OpenBookQA (Mihaylov et al., 2018) is a 4-way multiple choice QA dataset that requires reasoning with elementary science knowledge, containing 5,957 questions.
@@
MedQA-USMLE (Jin et al., 2020) is a 4-way multiple choice QA task that requires biomedical and clinical knowledge. The questions are originally from practice tests for the United States Medical License Exams (USMLE). The dataset contains 12,723 questions.
B.2. Graph-Text Pair Classification
@@
To evaluate Janus on graph-text pair classification, we curated a dataset based on WebNLG (Gardent et al., 2017). The original WebNLG dataset is used to assess models’ ability in text-to-graph and graph-to-text generation. Concretely, each data contains a graph-text pair where the graph is a set of triples from DBpedia and the text is the corresponding description of the triples. For example, the graph data are ”(John_E_Blaha birthDate 1942_08_26), (John_E_Blaha birthPlace San_Antonio), (John_E_Blaha occupation Fighter_pilot)”, while the corresponding text is ”John E Blaha, born in San Antonio on 1942-08-26, worked as a fighter pilot.”
@@
B.3. Retrieval
SciFact (Wadden et al., 2020) is a scientific fact-checking dataset aimed at verifying scientific claims using relevant research papers. It contains 920 training queries, and 300 test queries, with a corpus of 5,183 documents. In our experiment, we split 100 samples from the training set to serve as validation queries, ensuring that the remaining training data is used for model training while still providing a separate validation set for tuning and evaluation.
FiQA (Maia et al., 2018) is a financial-domain retrieval dataset designed to address complex financial question answering and information retrieval tasks. It contains 14,166 training queries, 500 development queries, and 648 test queries, with a total of 57,638 documents in the corpus.
Appendix C Methodology Details
C.1. Additional Framework Details
Janus utilizes a dual-view architecture for graph-text alignment, incorporating two parallel branches. The first branch processes the original graph together with the text, while the second branch processes a textual description of the graph along with the input text. In our experiments, both branches are employed to generate embeddings that are used for downstream tasks, and both contribute to the task-specific loss $\mathcal{L}^{\rm(task)}$. This dual-branch approach has several advantages: it allows the model to learn complementary information from both the structured graph and its textual description, enhancing the overall representation. During the evaluation, however, we simplify the process by using only the graph+text branch, and it has shown to provide sufficient performance for the downstream tasks, eliminating the need for the graph description branch.
C.2. Training Objective

### v2

@@ 4.3 Retrieval Performance
BERT-Base Devlin et al. (2019) 13.3 2.2
RoBERTa-Large Liu et al. (2019) 43.3 20.4
E5-Small Wang et al. (2022) 65.6 34.8
E5-Base Wang et al. (2022) 73.1 36.4
E5-Large Wang et al. (2022) 72.6 38.6
GTR-XXL Ni et al. (2022a) 66.2 46.7
SGPT Muennighoff (2022) 74.7 37.2
E5-Mistral, 0-shot Wang et al. (2023) 76.1 53.5
E5-Mistral, LoRA Wang et al. (2023) 75.9 53.2
@@
QA-GNN Yasunaga et al. (2021) 41.4 19.5
GreaseLM Zhang et al. (2022) 48.9 29.3
GT2Vec, LP (Ours) 82.9 54.1
GT2Vec, LoRA (Ours) 80.8 56.2
@@
Figure 3: (a) The effect of LLM backbone choice on accuracy for the CommonsenseQA dataset. The figure shows three series: E5, LLaMA-3, and LLaMA-2, along with a single Mistral-7B model. (b) The effect of graph encoder depth (number of GNN layers) on test accuracy for CommonsenseQA and OpenBookQA datasets. The shaded areas represent the standard deviation, indicating the variance in performance across different trials. (c) Graph-text embedding distance (red dashed) and dev accuracy (purple solid) on CommonsenseQA across training epochs.
As shown in Table 4, GT2Vec achieves the best performance on both datasets. A closer examination of the LM + KG baselines reveals interesting trends. Both models use RoBERTa-Large in their architectures. From the results, GreaseLM outperforms the RoBERTa-Large model by a margin of 5.6% on SciFact and 8.9% on FiQA, which aligns with our earlier findings that graph context can be useful for enhancing retrieval performance. However, QA-GNN shows a performance degradation compared to RoBERTa-Large, particularly on FiQA (19.5 vs. 20.4). The reason for this drop may be attributed to QA-GNN’s use of a simple MLP for combining the graph and text embeddings. This weaker integration mechanism is likely insufficient for fully leveraging the graph context, resulting in suboptimal performance. In contrast, GreaseLM employs a shallow transformer, which offers better capacity for fusing multimodal information, leading to moderate gains.
4.4 Ablation Studies
To better understand the contributions of different components in GT2Vec, we perform an ablation study on the CommonsenseQA and OpenBookQA datasets under the linear probe setting. More results can be found in Appendix E.2.
@@ 4.4 Ablation Studies
Effect of LLM Backbone Choice. GT2Vec exhibits flexibility in adopting various LLMs as its backbone. To evaluate the impact of different LLM backbones on performance, we compare four series of models: the E5 series Wang et al. (2022, 2023), LLaMA-2 Touvron et al. (2023), LLaMA-3 Dubey et al. (2024), and Mistral Jiang et al. (2023), as illustrated in Figure 3(a). We first observe that increasing the model size within each series consistently enhances the performance. This trend highlights that larger models, with more parameters, are typically able to better capture complex relationships in multi-modal data, leading to higher performance. Additionally, LLaMA-3 significantly improves over LLaMA-2. Specifically, the performance of LLaMA-3-3B matches that of LLaMA-2-7B and LLaMA-3-8B achieves results comparable to LLaMA-2-13B, indicating that the new LLaMA-3 architecture brings considerable advancements compared to its predecessor LLaMA-2.
@@
Effect of Graph Encoder Depth. We evaluate the effect of varying the number of GAT layers in the graph encoder on the performance of GT2Vec using the CommonsenseQA and OpenBookQA datasets. As shown in Figure 3(b), both datasets exhibit slight fluctuations in test accuracy as the number of GAT layers increases from 3 to 7. This suggests that the method is quite robust to changes in GAT depth, with only small variations in performance.
Effect of Graph-Text Alignment. We further investigate the effectiveness of graph-text alignment by studying how the alignment evolves during training. Specifically, we calculate the mean Euclidean distance between the normalized graph and text embeddings on the dev set (shown as the red dashed line) and compare it to the corresponding development accuracy (purple solid line) across different training epochs, as illustrated in Figure 3(c). This figure shows a clear inverse relationship between the two curves: as the graph-text distance decreases, accuracy improves. Notably, at epoch 2, the graph-text distance reaches its minimum, while the dev accuracy peaks. This trend suggests that better alignment between graph and text embeddings contributes directly to improved model performance, highlighting the effectiveness of our graph-text alignment strategy in GT2Vec.
5 Conclusion
@@
In this paper, we introduce GT2Vec, a simple yet effective framework designed to integrate graph and text data using a novel alignment strategy and LLMs. We have demonstrated that GT2Vec enhances the semantic coherence between these two modalities, resulting in significantly improved performance on several NLP tasks and datasets. By aligning graph embeddings directly with text embeddings, GT2Vec ensures a deeper integration of structured and unstructured data.
Limitations
While our proposed GT2Vec achieves strong performance across multiple tasks, it has certain limitations. First, our approach does not explicitly handle additional modalities such as images, which could provide richer contextual signals. Exploring effective ways to incorporate multi-modal data remains an important direction for future research. Second, although our model effectively aligns graph and text representations, there may be further opportunities to refine this alignment for better generalization across diverse tasks and domains.
@@ Appendix A Ethical Use of Data and Informed Consent
Appendix B Related Work
Graphs and Language Models for Multi-Modal Embedding Tasks. Integration of graph data with language models has been an evolving field, aiming to combine structured graph data with unstructured textual information for enhanced data analysis. Historically, early attempts in this area employed an MLP or a shallow transformer to merge the information from both modalities Lin et al. (2019b); Yasunaga et al. (2021); Zhang et al. (2022); Feng et al. (2020), which may not fully exploit the rich contextual information from text and graph data.
The advent of LLMs has brought a transformative shift to the NLP field. LLMs, with their extensive pre-training on diverse corpora, offer unprecedented capabilities for deep semantic understanding and reasoining Zhao et al. (2023b); Gao et al. (2023). Recent research has begun to explore the potential of LLMs for enhancing multimodal embedding tasks of graphs and text Huang et al. (2024); Tian et al. (2024); Zhao et al. (2023a); He et al. (2024a); Zhu et al. (2024c). These studies primarily focus on augmenting the graph representation capabilities within the LLM framework. While these efforts mark significant advancements, they predominantly concentrate on graph representation learning rather than direct NLP tasks. Additionally, these methods do not explicitly align the semantic spaces of graphs and text. In contrast, our approach not only integrates LLMs for handling complex multimodal inputs but also introduces an explicit alignment mechanism between graph and text embeddings. This alignment is crucial for enhancing the semantic integration and boosting performance across diverse NLP tasks by capturing complex intermodal relationships.
There are recent studies that have explored using LLMs and graph data for generative tasks beyond embedding tasks He et al. (2024b); Perozzi et al. (2024); Zhang et al. (2024); Chai et al. (2023); Hu et al. (2024); Jin et al. (2024); Zhu et al. (2024b). Our work, however, concentrates on embedding tasks, which is an orthogonal direction. Unlike generative tasks, which primarily focus on creating new content via next-token prediction, embedding tasks are aimed at developing rich, informative representations that can be used directly to enhance performance in downstream applications such as classification and retrieval.
Contrastive Learning. Contrastive learning has gained significant attention in recent years as a method to learn representations by maximizing agreement between positive pairs while pushing negative pairs apart in the embedding space He et al. (2020); Chen et al. (2020); Lin et al. ; Lin et al. (2024); Yu et al. (2022). Pioneering works like SimCLR Chen et al. (2020) and MoCo He et al. (2020) have demonstrated the effectiveness of contrastive learning in the visual domain. Similar approaches have been applied to graphs Yu et al. (2022); You et al. (2020); Fang et al. (2024) and language models Iter et al. (2020b); Fang et al. (2020); Giorgi et al. (2021), but these methods typically operate within a single modality (e.g., graphs or text). In contrast, our approach introduces contrastive learning to align the graph embedding space with the text embedding space. By transforming graphs into textual descriptions and applying contrastive learning, we ensure that embeddings from both modalities align closely, enhancing the model’s ability to perform tasks that require both graph-based reasoning and text comprehension.
Appendix C Dataset Details
C.1 KG-Contextulized QA
@@
CommonsenseQA Talmor et al. (2019) is a 5-way multiple-choice QA dataset focused on applying commonsense knowledge in answering questions. It includes 12,102 questions, each with one correct answer and four distractor answers.
@@
OpenBookQA Mihaylov et al. (2018) is a 4-way multiple choice QA dataset that requires reasoning with elementary science knowledge, containing 5,957 questions.
@@
MedQA-USMLE Jin et al. (2020) is a 4-way multiple choice QA task that requires biomedical and clinical knowledge. The questions are originally from practice tests for the United States Medical License Exams (USMLE). The dataset contains 12,723 questions.
For each question in the QA datasets, a subgraph context extracted from a KG is utilized to provide additional contextual information, following Yasunaga et al. Yasunaga et al. (2021). Specifically, for CommonsenseQA and OpeBookQA, we used the ConceptNet Speer et al. (2017), which contains 799,273 nodes and 2,487,810 edges. Node embeddings are initialized by Roberta-Large Liu et al. (2019) and kept frozen during the training process, following Yasunaga et al. (2021). The query node embeddings mentioned in Section 3.1 are calculated by the LLM backbone. For MedQA-USMLE dataset, we used the UMLS knowledge graph used in Yasunaga et al. (2021), which contains 9,958 nodes and 44,561 edges. Node embeddings are initialized using SapBERT Liu et al. (2021), following Yasunaga et al. (2021).
C.2 Graph-Text Pair Classification
@@
To evaluate GT2Vec on graph-text pair classification, we curated a dataset based on WebNLG Gardent et al. (2017). The original WebNLG dataset is used to assess models’ ability in text-to-graph and graph-to-text generation. Concretely, each data contains a graph-text pair where the graph is a set of triples from DBpedia and the text is the corresponding description of the triples. For example, the graph data are "(John_E_Blaha birthDate 1942_08_26), (John_E_Blaha birthPlace San_Antonio), (John_E_Blaha occupation Fighter_pilot)", while the corresponding text is "John E Blaha, born in San Antonio on 1942-08-26, worked as a fighter pilot."
@@
C.3 Retrieval
SciFact Wadden et al. (2020) is a scientific fact-checking dataset aimed at verifying scientific claims using relevant research papers. It contains 920 training queries, and 300 test queries, with a corpus of 5,183 documents. In our experiment, we split 100 samples from the training set to serve as validation queries, ensuring that the remaining training data is used for model training while still providing a separate validation set for tuning and evaluation.
FiQA Maia et al. (2018) is a financial-domain retrieval dataset designed to address complex financial question answering and information retrieval tasks. It contains 14,166 training queries, 500 development queries, and 648 test queries, with a total of 57,638 documents in the corpus.
Appendix D Methodology Details
D.1 Additional Framework Details
GT2Vec utilizes a dual-view architecture for graph-text alignment, incorporating two parallel branches. The first branch processes the original graph together with the text, while the second branch processes a textual description of the graph along with the input text. In our experiments, both branches are employed to generate embeddings that are used for downstream tasks, and both contribute to the task-specific loss $\mathcal{L}^{\rm(task)}$. This dual-branch approach has several advantages: it allows the model to learn complementary information from both the structured graph and its textual description, enhancing the overall representation. During the evaluation, however, we simplify the process by using only the graph+text branch, and it has shown to provide sufficient performance for the downstream tasks, eliminating the need for the graph description branch.
D.2 Training Objective